OpenStudy (anonymous):

Help me find the differential dy given y=(√3x^2-4), then dy = a. 6x(√3x^2-4)dx b. (3x)/(√3x^2-4)dx c. (6x)/(√3x^2-4) d. 2(√3x^2-4)/(3x)dx

OpenStudy (anonymous):

is that seperate questions? im not understanding it..

OpenStudy (anonymous):

it goes from a , b , c , d? :/

OpenStudy (alexwee123):

@Omniscience i think they are answer choices

OpenStudy (anonymous):

multiple choice? the asker should confirm..

OpenStudy (anonymous):

yes they are. bleh

OpenStudy (anonymous):

can you please write it properly? :)

OpenStudy (anonymous):

yea just give me a sec

OpenStudy (alexwee123):

hint* use chain rule

OpenStudy (anonymous):

a) \[6x \sqrt{3x^2-4}dx\] b)\[3x/\sqrt{3x^2-4}dx\] c)\[6x/\sqrt{3x^2-4}dx\] d)\[2\sqrt{3x^2-4}/3x dx\]

OpenStudy (alexwee123):

F'(x) = f '(g(x)) g '(x) => chain rule so if we apply it....